# Designing ReLU Generative Networks to Enumerate Trees with a Given Tree Edit Distance

### v1

@@ 6 TEd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TEd-generative ReLU to be a ReLU neural network such that each Euler string over $\Sigma$ with edit distance at most $2d$ from $E(T)$ due to deletion, substitution and insertion operations can be obtained by appropriately choosing an input $x=x_{1},x_{2},\ldots,x_{7d}$ of $7d$ nodes with $x_{j}\in[0,1)$, where $x_{j}$ is of the form $\cdot\Delta, ii$ is an integer and $\Delta$ is a small constant. The input $x_{1},x_{2},\ldots,x_{d}$ (resp., $x_{d+1},x_{d+2},\ldots,x_{3d}$ and $x_{3d+1},x_{3d+2},\ldots,x_{7d}$) represents the nodes for deletion (resp., substitution and insertion) operations. As a preprocessing step, the random inputs $x_{j}$ for $1\leq j\leq 2d$ and $3d+1\leq j\leq 6d$ (resp., $2d+1\leq j\leq 3d$ and $6d+1\leq j\leq 7d$) are converted into integers $i\in\{0,\ldots,n\}$ (resp., $\ell\in\Sigma$) if $x_{j}\in((i-1)/n,i/n]$ (resp., $x_{\ell}\in[(\ell-1)/m,\ell/m]$ for $\ell=1$, $((\ell-1)/m,\ell/m]$ otherwise). For example, when $n=5$ and $m=10$, the conversion table is given in Table 7 in Appendix 9. To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, where $B\gg\max(m,n)$. The network outputs $2n+2d$ nodes $y=y_{1},\ldots,y_{2n+2d}$ from which the desired string $E(U)$ can be obtained by trimming all $B$s from the start and end. More precisely, if $d_{1}$ (resp., $d_{2}$) deletion (resp., insertion) operations are performed, then $2d-2d_{1}$ (resp., $2d_{2}$}) number of $B$s will be trimmed from the end (resp., start) of the output $y$ as shown in Fig. 9 in Appendix 9.
Figure 9: An illustration of a TEd-generative ReLU for $d=3$, with the input layer x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02x=0.3,0,0.38,0,0.46,0.55,0,0.6,0.88,0.66,0.75,0,0.55,0.87,0.03,0.02, $0.45,0.09,0,0.7,0.5$, padded Euler string $3,2,12,2,4,14,12,4,14,13,B,B,B,B,B,B$ and the output layer $B,B,B,B,5,3,2,6,16,12,4,14,13,15,B,B,B,B$ with $d_{1}=d_{2}=1$. By trimming $B$s, we can get the resultant string $E(U)=5,3,2,6,16,12,4,14,13,15$ obtained by deleting, substituting and inserting the indicated nodes.
@@ 7 Computational Experiments
We implemented the proposed networks on a machine with an AMD Ryzen 7 4800H, Radeon Graphics processor (2.90 GHz), 16 GB of RAM, and Windows 11 Pro using Python version 3.11.6.
@@
For each tree $T_{i}$, $i=1,2,3,4$ illustrated in Figs. 10(a)-(d) with $8,10,6,11$ nodes, labels from $\Sigma=\{1,2,\ldots,10\}$, and $d=2,2,3,2$, resp., we generated all non-isomorphic trees $U$ that have tree edit distance exactly $d$ by using the proposed TId-generative ReLU networks. The newly inserted nodes in $T_{1}$ are 7, 7, in $T_{2}$ are 9, 9, in $T_{3}$ are 8, 8, 8, and in $T_{4}$ are 2, 2. The computational results such as the number of nodes in each layer of the constructed TId network along with the number of generated trees for each input tree by the network are provided in Table 4. A summary of these computational results is given below. For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TId-generative ReLU network for $T_{i}$. From these experiments, we have $(57,32876,6,576.77,11903)^{\rm I}_{1}$, $(57,57324,6,1005.68,24841)^{\rm I}_{2}$, $(57,20556,9,360.63,4500)^{\rm I}_{3}$, $(57,73064,6,1281.82,33860)^{\rm I}_{4}$. These values are illustrated in Fig. 11(a) for each tree $T_{i}$, $i=1,2,3,4$. Observe that the depth remains fixed across all trees, confirming Theorem 3. Moreover, the widest layers are significantly larger than the others; for example, in $T_{4}$ with 11 nodes, the widest layer contains 33860 nodes which is over 26.41 times the average layer size (1281.82). Additionally, the total number of nodes grows faster than the tree size, e.g., while 11 nodes are roughly twice 6, the corresponding total node count (73064) is 3.56 times that for 6 nodes (20556).
Figure 11: Information of the architectures of the proposed ReLU-networks for insertion and unified cases and their comparison: (a) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm I}_{i}$ of the proposed ReLU-networks for insertion case for each tree $T_{i}$, $i=1,2,3,4$; (b) Information of $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ of the proposed ReLU-networks for unified case for each tree $T_{i}$, $i=1,2,3,4,5$; (c) A comparison of the information of $(\#\rm{TN},\rm{AvgN},\#\rm{MaxN})_{i}$ of the proposed ReLU-networks for insertion and unified cases for each tree $T_{i}$, $i=1,2,3,4$.
@@
Input trees Size of each hidden layer of TId-generative ReLU Number of generated trees
@@
The computational results of these experiments are given in Table 5 which are summarized below. For $T_{i}$, let $(\#\rm{L},\#\rm{TN},\rm{MinN},\rm{AvgN},\#\rm{MaxN})^{\rm E}_{i}$ denote the sequence of number of hidden layers, total number of hidden nodes, minimum number of hidden nodes, average number of hidden nodes, and maximum number of hidden nodes, resp., in the TEd-generative ReLU network for $T_{i}$. For the computational results listed in Table 5 we have $(108,50360,14,466.30,11917)^{\rm E}_{1}$, $(108,82060,14,759.81$, $24859)^{\rm E}_{2}$, $(108,35803,19,331.51,4510)^{\rm E}_{3}$, $(108,101930,14$, $943.80,33880)^{\rm E}_{4}$, $(108,493790,14,4572.13,263350)^{\rm E}_{5}$ for TEd-generative ReLU networks. These values are illustrated in Fig. 11(b) for each tree $T_{i}$, $i=1,2,3,4,5$.
Moreover a comparison of the TId and TEd is given in Fig. 11(c). Observe that TEd networks are deeper and, on average, narrower than TId networks for the same tree. Similarly, TEd networks exhibit larger minimum widths compared to TId networks, while the maximum layer widths are nearly identical in both cases. These findings indicate that the insertion operation contributes the most to the overall size of the unified networks, which are composed of substitution, deletion, and insertion components. This observation also supports Theorems 3 and 4.
@@
Input trees Size of each hidden layer of TEd-generative ReLU Number of generated trees
@@ 8 Conclusion
We study the existence of ReLU-based generative networks for producing trees similar to a given tree with respect to the tree edit distance. Our approach transforms a rooted, ordered, and vertex-labeled tree into a rooted, ordered, and edge-labeled directed tree. This directed tree is then encoded as an Euler string, which serves as both the input and output of the ReLU generative networks. First, we proved that there exists a ReLU network of size $\mathcal{O}(dn)$ and constant depth that can identify the labels of $d$ inward edges in the Euler string. Furthermore, we showed that the outward edges corresponding to these $d$ inward edges can be identified using a ReLU network of size $\mathcal{O}(dn^{2})$ and constant depth. Building on these results, we demonstrated that all similar trees generated through substitution (resp., deletion and insertion) operations can be constructed by ReLU networks of size $\mathcal{O}(dn^{2})$ (resp., $\mathcal{O}(n^{2})$ and $\mathcal{O}(n^{3})$), all with constant depth. Finally, we proved that there exists a ReLU network of size $\mathcal{O}(n^{3})$ and constant depth that can generate all trees with a distance at most $d$ to the original tree under combined substitution, deletion, and insertion operations. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for efficient tree-structured data generation.

### v3

@@ 6 TEd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TEd-generative ReLU to be a ReLU neural network such that each Euler string over $\Sigma$ with edit distance at most $2d$ from $E(T)$ due to deletion, substitution and insertion operations can be obtained by appropriately choosing an input $x=x_{1},x_{2},\ldots,x_{7d}$ of $7d$ nodes with $x_{j}\in[0,1)$, where $x_{j}$ is of the form $\cdot\Delta, ii$ is an integer and $\Delta$ is a small constant to set the number of digits to be considered after decimal in $x_{j}$. For example, if $\Delta=0.01$, then $x_{j}$ can be any number in [0,1) with two decimal places, i.e., $x_{j}$ cannot take the value 0.011. . The input $x_{1},x_{2},\ldots,x_{d}$ (resp., $x_{d+1},x_{d+2},\ldots,x_{3d}$ and $x_{3d+1},x_{3d+2},\ldots,x_{7d}$) represents the nodes for deletion (resp., substitution and insertion) operations. As a preprocessing step, the random inputs $x_{j}$ for $1\leq j\leq 2d$ and $3d+1\leq j\leq 6d$ (resp., $2d+1\leq j\leq 3d$ and $6d+1\leq j\leq 7d$) are converted into integers $i\in\{0,\ldots,n\}$ (resp., $\ell\in\Sigma$) if $x_{j}\in((i-1)/n,i/n]$ (resp., $x_{\ell}\in[(\ell-1)/m,\ell/m]$ for $\ell=1$, $((\ell-1)/m,\ell/m]$ otherwise). For example, when $n=5$ and $m=10$, the conversion table is given in Table 7 in Appendix 9. To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, where $B\gg\max(m,n)$. The network outputs $2n+2d$ nodes $y=y_{1},\ldots,y_{2n+2d}$ from which the desired string $E(U)$ can be obtained by trimming all $B$s from the start and end. More precisely, if $d_{1}$ (resp., $d_{2}$) deletion (resp., insertion) operations are performed, then $2d-2d_{1}$ (resp., $2d_{2}$}) number of $B$s will be trimmed from the end (resp., start) of the output $y$ as shown in Fig. 9 in Appendix 9.
Figure 9: An illustration of a TEd-generative ReLU for $d=3$, with the input layer x=x=0.3, 0, 0.38, 0, 0.46, 0.55, 0, 0.6, 0.88, 0.66, 0.75, 0, 0.55, 0.87, 0.03, 0.02, $0.45,0.09,0,0.7,0.5$, the integer conversion of $x$ in 2, 0, 2, 0, 3, 3, 1, 6, 9, 4, 4, 0, 3, 5, 1, 1, 3, 1, 1, 7, 5, padded Euler string $3,2,12,2,4,14,12,4,$ $14,13,B,B,B,B,B,B$ and the output layer $B,B,B,B,5,3,2,6,16,12,4,14,13,15,B,B,B,B$ with $d_{1}=d_{2}=1$. By trimming $B$s, we can get the resultant string $E(U)=5,3,2,6,16,12,4,14,13,15$ obtained by deleting, substituting and inserting the indicated nodes.
@@ 7 Computational Experiments
We implemented the proposed networks on a machine with an AMD Ryzen 7 4800H, Radeon Graphics processor (2.90 GHz), 16 GB of RAM, and Windows 11 Pro using Python version 3.11.6. Note that the proposed $\mathrm{TE}_{d}$-generative ReLU network supports all three edit operations, deletion, substitution, and insertion, whereas the $\mathrm{TI}_{d}$-generative ReLU network supports only insertion operations; however, $\mathrm{TI}_{d}$ has the same order $\mathcal{O}(n^{3})$ as the corresponding $\mathrm{TE}_{d}$ Therefore, for analysis and comparison, we generated trees using the $\mathrm{TI}_{d}$ and $\mathrm{TE}_{d}$ networks for four trees $T_{i}$, $i=1,2,3,4$, shown in Fig. 10, with $8,10,6,$ and $11$ nodes, respectively, labels from $\Sigma=\{1,2,\ldots,10\}$, and distance bounds $d=2,2,3,$ and $2$.
@@
For each tree $T_{i}$, $i=1,2,3,4$ illustrated in Figs. 10(a)-(d), we supplied the proposed $\mathrm{TI}_{d}$-generative ReLU network with a collection of input sequences $x$ that can cover all possible insertion operations to generate all possible similar trees with distance exactly $d$ due to the insertion operations. Duplicate outputs were removed in a post-processing step, yielding the set of distinct trees $U$ whose tree edit distance from $T_{i}$ is exactly $d$. In these experiments, two (resp., two; three; and two) new nodes were inserted with labels 7, 7 (resp., 9, 9; 8, 8, 8; and 2, 2) in $T_{1}$ (resp., $T_{2}$; $T_{3}$; and $T_{4}$) to generate similar trees of edit distance exactly $2$ (resp., 2; 3; and 2) from $T_{1}$ (resp., $T_{2}$; $T_{3}$; and $T_{4}$) using the proposed $\mathrm{TI}_{d}$ network. The computational results such as the number of nodes in each layer of the constructed TId network along with the number of all distinct trees obtained after removing duplication from the similar trees generated by the proposed network for each input tree are provided in Table 4. A summary of these computational results is given below. For $T_{i}$, let $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},RT)^{\rm I}$i denote the sequence of number of hidden layers in TId, total number of hidden nodes in TId, minimum number of hidden nodes in TId, average number of hidden nodes in TId, maximum number of hidden nodes in TId, and running time (sec.) to generate one similar tree using TId-generative ReLU network for $T_{i}$, respectively. From these experiments, we have $(57,32876,6,576.77,11903,17.65)^{\rm I}_{1}$, $(57,57324,6,1005.68,24841,46.23)^{\rm I}_{2}$, $(57,20556,9,360.63,$ $4500,7.11)^{\rm I}_{3}$, $(57,73064,6,1281.82,33860,69.84)^{\rm I}_{4}$. These values are illustrated in Fig. 11(a) for each tree $T_{i}$, $i=1,2,3,4$. Observe that the depth remains fixed across all trees, confirming Theorem 3. Moreover, the widest layers are significantly larger than the others; for example, in $T_{4}$ with 11 nodes, the widest layer contains 33860 nodes which is over 26.41 times the average layer size (1281.82). Additionally, the total number of nodes grows faster than the tree size, e.g., while the number of nodes (11) in $T_{4}$ are roughly 1.8 times the number of nodes (6) in $T_{3}$, the corresponding total node count (73064) in TId for $T_{4}$ is 3.56 times the node count (20556) in TId for $T_{3}$. As a result, the running time also increases faster than the tree size.
Figure 11: Information of the architectures of the proposed ReLU-networks for insertion and unified cases and their comparison: (a) Information of $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},$ $\rm{MaxN},\rm{RT})^{\rm I}$i of the proposed ReLU-networks for insertion case for each tree $T_{i}$, $i=1,2,3,4$; (b) Information of $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},{\rm RT})^{\rm E}$i of the proposed ReLU-networks for unified case for each tree $T_{i}$, $i=1,2,3,4,5$. The ranges of L, MinN, AvgN, and RT are much smaller than those of TN and MaxN, which causes their plots to appear compressed and overlap when shown on the same scale; (c) A comparison of the information of $(\rm{TN},\rm{AvgN},\rm{MaxN},\rm{RT})$i of the proposed ReLU-networks for insertion and unified cases for each tree $T_{i}$, $i=1,2,3,4$.
@@
Input trees Size of each hidden layer of TId-generative ReLU Number of all distinct similar trees obtained by TId
@@
The detailed computational results of these experiments are given in Table 5 which are summarized below. For $T_{i}$, let $(\rm{L},\rm{TN},\rm{MinN},\rm{AvgN},\rm{MaxN},\rm{RT})^{\rm E}$i denote the sequence of number of hidden layers in TEd, total number of hidden nodes, in TEd, minimum number of hidden nodes in TEd, average number of hidden nodes in TEd, maximum number of hidden nodes in TEd, and running time (sec.) to generate one similar tree using TEd-generative ReLU network for $T_{i}$, respectively. We have $(108,50360,14,466.30,11917,25.59)^{\rm E}_{1}$, $(108,82060,14,759.81$, $24859,55.23)^{\rm E}_{2}$, $(108,35803,19,$ $331.51,4510,11.36)^{\rm E}_{3}$, $(108,101930,14$, $943.80,33880,$ $85.04)^{\rm E}_{4}$, $(108,493790,14,$ $4572.13,263350,1581.17)^{\rm E}_{5}$ for TEd-generative ReLU networks. These values are illustrated in Fig. 11(b) for each tree $T_{i}$, $i=1,2,3,4,5$. Moreover a comparison of the TId and TEd is given in Fig. 11(c). Observe that TEd networks are deeper and, on average, narrower than TId networks for the same tree. Similarly, TEd networks exhibit larger minimum widths compared to TId networks, while the maximum layer widths are nearly identical in both cases. These findings indicate that the insertion operation contributes the most to the overall size of the unified networks, which are composed of substitution, deletion, and insertion components. This observation also supports Theorems 3 and 4.
@@
Input trees Size of each hidden layer of TEd-generative ReLU Number of all distinct similar trees obtained by TEd
@@ 8 Conclusion
We study the existence of ReLU-based generative networks for producing trees similar to a given tree with respect to the tree edit distance. Our approach transforms a rooted, ordered, and vertex-labeled tree into a rooted, ordered, and edge-labeled directed tree. This directed tree is then encoded as an Euler string, which serves as both the input and output of the ReLU generative networks. First, we proved that there exists a ReLU network of size $\mathcal{O}(dn)$ and constant depth that can identify the labels of $d$ inward edges in the Euler string. Furthermore, we showed that the outward edges corresponding to these $d$ inward edges can be identified using a ReLU network of size $\mathcal{O}(dn^{2})$ and constant depth. Building on these results, we demonstrated that all similar trees generated through substitution (resp., deletion and insertion) operations can be constructed by ReLU networks of size $\mathcal{O}(dn^{2})$ (resp., $\mathcal{O}(n^{2})$ and $\mathcal{O}(n^{3})$), all with constant depth. Finally, we proved that there exists a ReLU network of size $\mathcal{O}(n^{3})$ and constant depth capable of generating any tree within distance $d$ from the original tree under combined substitution, deletion, and insertion operations. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for efficient tree-structured data generation.